OpenStudy (anonymous):

A line passes through (1, -4( and has a slope of 3. What is an equation of the line in point-slope form?

OpenStudy (nubeer):

m=3 , (1,-4) y-y1=m(x-x1) just plug in the values u will get the equation.

OpenStudy (anonymous):

So confusing

OpenStudy (nubeer):

hmm which part?

OpenStudy (anonymous):

Just everything

OpenStudy (anonymous):

I learn backwards, I usually get the answer and go back to see how you got it. Iv always learned better like that.

OpenStudy (nubeer):

y-y1=m(x-x1) this is the general equation in which you have to plug in values m=slope= 3 (1,-4) x1=1, y1 = -4 now just put these in the general equation.